OpenStudy (anonymous):

Jack usually mows his lawn in 5 hrs, marilyn can mow the same yard in 4 hrs. How much time would it take for them to mow the lawn together?

OpenStudy (anonymous):

x/5 + x/4 = 1 20( x/5 +x/4=1) 4x+5x=20 9x=20 x=20/9 = 2.4444444444 hours

OpenStudy (anonymous):

Jack will mow 1/4 per hour. Marilyn will mow 1/6 per hour. 1/4 + 1/6 = 3/12 + 2/12 = 5/12 Together they will mow 5/12 per hour. Total hours x mowed per hour = 1 Hours x 5/12 = 1 Hours = 1/(5/12) Hours = 2.4

OpenStudy (anonymous):

1/5 not 1/6 but that way will work very well too

OpenStudy (anonymous):

sorry my bad.

OpenStudy (anonymous):

BOTH OF YOU WERE WRONG BUT THANK YOU FOR YOUR EFFORTS

OpenStudy (anonymous):

what is the answer plshlpme

OpenStudy (anonymous):

IT WAS 2 2/9

OpenStudy (anonymous):

which is 20/9. I just converted to a decimal instead of a mixed number, sorry

OpenStudy (anonymous):

ITS OK! :) THANK YOU FOR TRYING.